Provide a border and title block on two 8.5 x 11 sheets in portai approximately equal rectangles. Provide each section with a catorientation, origin in its center. Use 0.1 inch units (10 units per inch). Plotesiarn form a vector in each region. The length of the line is its magnitud conet th flsitem having the vector is given along with the pair of provided potsProviede protractor to measure the 1.) magnitude and 2.) direction (deg 3.) X and 4.)Y components. Direction is always counterclockwise f each vector as magnitudes and directions with calculations using the Pythagorearm , axis. Chek th the vector functions. Neatly show all work in each square. Carte Divide each into 4 coordinate system having the e general direction for each efrow he rees ads.) Use your scale and s) of each vector as well as the vector wise from o,X axis. Check the vector and the sin, cos, or tan 1) (0.0,0.0) &(2,0.5 Up, right Down, right 3.)0.0,0.0) & (-1.5, 1) Up, left 4) (0.0,0.0) & (0.4, 1.5) Up, right Down, left Up, left Down, right Up, left 2) 3J 4) 025 Border 3 Letters 16
media%2F0d2%2F0d21d7eb-0a43-49ae-8f08-90
media%2F225%2F225b41ba-1c62-412d-86b8-c4
Answer #1

BLOCK:

SLOPED SLAB:

Percent Grade of Slope = Rise/Run = 4.5/9 = 0.5 = 50%